a town has two fire engines operating independently. The probability that a specific fire engine is available when needed is 0.96
a) what is the probability that neither is available when a needed?
b)what is the probability that a fire engine is available when needed?

Answer by math_helper(2461)

P(a specific fire engine is NOT available) = 1 - P(it is available) = 1-0.96 = 0.04
a) 0.04*0.04 = 0.0016
b) 1-0.0016 = 0.9984